Problem: A rock is 100m above the ground. It is given an initial velocity of -1m/s (downward). The acceleration due to gravity is -9.8 m/s. How far above the ground is it after three seconds?
Answer #1

displacement travelled by the rock in hmgiven time

d = ut + 0.5 gt^2

d = 1 * 3 + 0.5* 9.8* 3^2

d = 47.1 m

Distance above ground = 100 - 47.1 = 52.9 m

==========

Comment in case any doubt, will reply for sure.. Goodluck